find the equation pf a line that is perpendicular to the line y=1/8x+9 and contains the point (-4,0)
y=

Answer by erica65404(394)

we know the equation of a line is y = mx + b.
The slope of y = 1/8x + 9 is 1/8.
A perpendicular line has slope that is the reciprical multiplied by -1.
So the new slope of the perpendicular equation is -8.
y = -8x + b
Now to find b we will plug in the x and y value of the point given into the equation.
0 = -8(-4) + b
0 = 32 + b
-32 = b
Now to finish the equation plug b into the equation.
y = -8x - 32
This is the answer.
-------------------------------------------------------------------------------
Go to mathhelp.createaforum.com for more help by me or email me at ericahigley@yahoo.com.